Subject: 3.5-m Users Committee minutes 2/25/08
From: Bruce Gillespie
Submitted: Wed, 5 Mar 2008 12:23:05 -0700
Message number: 1092
(previous: 1091,
next: 1093
up: Index)
APO 3.5-m Users Committee Phonecon, 2/25/08
Attending: Suzanne Hawley, Bruce Gillespie, Scott Anderson, John
Bally, Remy Indebetouw, Al Harper, Jon Holtzman, Jon Fulbright, Ed
Turner (for Michael Strauss)
Absent: Michael Strauss, Mark Klaene
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Turner) - Ed sat in on the telecon for Michael, and he had
nothing new to report on user feedback or issues.
UVa (Indebetouw) - Remy had nothing new to report on user feedback or
issues.
UW (Anderson) - Scott reported that a user at UW, Rich Plotkin, has
noticed the NIC-FPS bright-target persistence issue. It has not
caused him problems, but he is generally working with fainter
objects. Plotkin has noticed that the persistence lasts for 10s of
minutes for 5 second exposures for targets brighter than H = 11
magnitudes.
CU (Bally) - John had nothing new to report on user feedback or
issues. The new, better shielded, fan-out board for the NIC-FPS
detector is being fabricated. Its installation in the instrument, and
the removal of the faulty F-P etalon for testing at CU, are expected
in April.
JHU (Fulbright) - Jon had nothing new to report on user feedback or
issues.
NMSU (Holtzman) - Jon had nothing new to report on user feedback or
issues.
UChicago (Harper) - Al had nothing new to report on user feedback or
issues.
**********************************
Discussion of telescope/instruments report - Gillespie, Hawley:
The report is given below. Bruce and Suzanne went over the high
points of the report; there was little substantive discussion.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 1/07/08
through 2/20/08
Mark Klaene & Bruce Gillespie
0) Overview
Operations continued largely uninterrupted. Snowfalls to date have
been light and have not caused any significant loss of observing time
from snow and ice buildup on the roof. The recent lack of
precipitation is continuing to worry local forest fire officials about
the fire risk this spring. The severe cold weather should be behind
us; we saw no low-temperature related failures or problems this year.
1) Telescope
There were no issues with the telescope during this period. We are
successfully running the new axis controller on the instrument rotator
during science observing.
2) Instruments
DIS operated smoothly throughout the period, with no reported
problems. SPIcam had a shutter problem traced to the shutter wheel
coming loose from the motor shaft. This particular shutter had been
reliable for over 5 years, and the shutter wheel has since been
reinstalled and secured in place. We noticed a burning smell in the
Echelle electronics rack--we couldn't pinpoint the exact source but it
is suspected to have been coming from one of the unused electronics
cards. A test of all Echelle normal functions showed that everything
is working as expected. We will perform a more thorough search during
the summer, but for now the odor has disappeared. We are currently
waiting for the new fanout board for NIC-FPS and then will schedule a
date to warm the instrument, install the fanout board, and remove the
etalon for testing back at CU. We will try to do this work during
unscheduled observing time, which will likely be during April due to
TripleSpec's arrival in March. The newly coated focal reducer for
Agile was tested for throughput; preliminary data reduction and
analysis are complete and the web documentation has been updated. The
visiting instruments (CorMASS, GFP, and APOLLO) were all operational.
We plan to remove CorMASS from the observing level floor on March 3rd
in preparation for TripleSpec, after which time CorMASS will not be
available for regularly scheduled observing.
3) CIF projects
We are progressing on the implementation of Agile field rotation and
its migration to the TR2 mirror cell port. Design drawings are going
to the shop, and parts have been ordered. The Critical Design Review
for the Tertiary support/actuator upgrade is scheduled for March 6 and
7th.
**********************************
ARC collegiality and sharing of data - Hawley:
For context, Suzanne reminded us that we developed a Target of
Opportunity policy a few years back, largely to handle GRBs, and the
policy contains language about data rights and sharing. Recently,
there were two sequential requests for Director's Discretionary time
asking to take ToO observations of a SNe event. The existing GRB
policy was largely designed for fast-attack preemptions, and the
slower response time needed for the SNe programs did not entirely fit
the GRB guidelines. In addition, the two SNe observing groups
consisted of several team members from non-ARC institutions, who were
in competition with each other. This resulted in resistance to
sharing the DD observing time and data. Suzanne wondered if there
needs to be a more general guideline for ToOs regarding shared
observing and data rights, in order to enhance both scientific
efficiency and ARC collegiality. Ed Turner, who was sitting in for
Michael, said that when the ARC Board of Governors originally
established DD time, it was envisaged as being administered entirely
at the Director's discretion. There were no guidelines established
other than that the Director would report on DD usage at the annual
Board meetings. He agreed that ToOs are thorny problems to deal with,
and it is probably impossible to share observing time with those
requiring fast-attack (i.e., hours, not days) preemptions of scheduled
programs. Since a wide variety of actions can be required to
accommodate any of a large diversity of possible DD requests, Ed felt
that it would be better not to have general policies, but rather to
give the Director broad latitude in what course of action to take. Al
Harper concurred that it would be best to leave it to the Director.
As for the collegiality issue, Ed said that it is perfectly legal (in
the ARC sense) that the partner institutions can do whatever they want
with the data taken at APO, although our community culture has been
largely generous about sharing data. However, when outside interests
are involved in data rights, issues will naturally arise. Al suggested
when there are data-sharing issues involving outside collaborations,
we should be tending towards collegiality whenever possible. John
Bally also said he supports this overall general approach.
Suzanne said she would consider these points, and invited the
committee members and others to send her any further opinions on the
matter.
**********************************
TripleSpec update - Hawley:
TripleSpec has undergone several cool-down cycles in the lab at UVa,
and except for some final focus tweaks, it appears close to being
ready for shipment to APO. Lab calibration data are looking good, and
there are a few remaining issues to be worked, such as finishing the
instrument control software and TUI interface, and instrumental heat
advection. Shipment to APO is targeted for the 2nd week in March, and
commissioning on the telescope the following week during engineering
time. Additional engineering time in Q2 will be allocated for
continued testing, and TripleSpec might be in a state where a small
number of shared-risk science observations may be possible later in
the spring. There is no scheduled TripleSpec science time in Q2
because of the natural uncertainties of how fast the commissioning
will proceed, but interested users who have a couple of targets are
invited to contact John Wilson (jcw6z at virginia.edu) to see if these
can be observed during the later phases of commissioning on a shared-
risk, best-effort basis. By early May, we hope to be able to decide
if TripleSpec will be offered for scheduled science time in Q3. There
is some user information on TripleSpec already linked to on our user
information webpages. Remy will check to see if these are as up-to-
date as possible.
**********************************
Newly trained observers protocol - Hawley:
Regarding the training policy for new users, Suzanne mentioned that
newly trained users are not automatically added to the list of trained
observers after they have completed their three-night orientation at
the observatory. To be added to the list of authorized remote
observers (i.e., trained), we ask that either the person's
institutional scheduler or Users Committee representative send an e-
mail to this effect to Suzanne, Mark, Bruce, and Gretchen so that
their name and contact information can be added to the trained user
list.
**********************************
NIC-FPS raw data curation policy - Hawley:
With typically short exposure times and multiple raw data frames
resulting from Fowler sampling, NIC-FPS can produce prodigious amounts
of raw data. Our current NIC-FPS data processing system returns
processed frames immediately to the user, but the raw-frame data are
kept on disk at APO, from where the user can download them at a later
time. Although we normally keep all data for the other instruments for
approximately 9 months, we are going to establish a practice that
purges raw NIC-FPS frames 90 days after the observations are taken.
If NIC-FPS users are interested in having their raw data frames, we
strongly suggest that they download these data not later than a few
weeks after they are taken. John Bally suggested implementing a
system where the raw NIC-FPS frames are time-stamped so that they can
be automatically deleted after exactly three months, and Suzanne
thought this would be possible. Remy said that this policy should be
written into the NIC-FPS user documentation, and Suzanne agreed.
**********************************
Telescope tracking while mirror covers closed - Gillespie/Hawley:
Some observers ask to do lamp calibrations with the primary mirror
covers closed while continuing to track their targets, in order to get
more calibration signal off the whitened surfaces on the mirror
covers, and to save time by not having to reacquire the target. By
design, the telescope is significantly out of balance when the mirror
covers are closed, and tracking the telescope in this condition puts
strain on the drive electronics and mechanical components. We are
establishing the standard practice that telescope tracking will always
be turned off when the primary mirror covers are in the closed position.
**********************************
ACTION ITEMS:
[open from previous months]:
ACTION: John Bally will write a section with figures on Fowler
sampling for inclusion in the NIC-FPS user documentation.
STATUS: Closed, perhaps. Cory has given Gabrelle new materials for
the NIC-FPS user information pages.
ACTION: User Committee members should find out if their NIC-FPS users
are having problems with residual images caused by bright targets.
STATUS: Closed. Cory has taken data on the persistence problem,
which will be written up as a guide in the NIC-FPS documentation.
John Bally said that clearing the bright-target residual images may
take hours and could be a scheduling issue. We will need to keep
reminding NIC-FPS users of this instrumental characteristic.
[new actions from this month]:
none
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 31 March, at 8:30
AM Pacific Time. The agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1092 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO